# A Longitudinal Study of Child Wellbeing Assessment via Online Interactions with a Social Robot

### v2

@@ 3 Research Questions and Contributions
Children differ widely in their attitudes towards technology which might affect their interaction with a social robot [64]. This is of specific importance when encouraging them to open up about their wellbeing. It has also been shown that wellbeing is influenced by multiple factors (family problems, financial issues, academic workload etc.) and is subject to change without any warning or discerning pattern [65]. Thus, we investigated whether and how online interaction (using Zoom or Skype) with a social robot can be used for the longitudinal assessment (three sessions over the course of 3-6 weeks) of mental wellbeing (RQ1). Longitudinal CRI studies have also observed that habituation with the robot leads to children losing interest in the interaction, causing less engagement and attention to tasks [40]. Thus, we investigated how children’s perception of the robot changed over the course of the longitudinal interaction (RQ2). Specifically, we also investigated whether remote interactions with the robot affect the children’s perception of the robot as compared to co-location (RQ2.1). Gender-based differences have also been observed with regard to participants’ attitudes towards mental health [13] and their HRI experience [50]. Thus, we also conducted a supplementary exploratory investigation into whether and how gender affects the robot-assisted longitudinal assessment of the mental wellbeing of children and their perception of the robot’s behaviour and capabilities (RQ3).
@@
Figure 2: (a) Researcher view of the experimental setup, (b) Participant view of the interaction.
4 The Study
@@
The experiment design for the present work was motivated by our prior investigation examining whether and how a humanoid robot can be used for the evaluation of mental wellbeing of children in a dyadic physical lab setting [4, 62]. In our in-lab study, we collected data from 41 children [^5] In order to balance out the gender of the participants, we collected data from about 13 more participants after the 28 participants mentioned in [4] (21 girls, 20 boys, mean age = $9.58y.o.$, SD = $1.45y.o.$). The experimental protocol consisted of the robot following a pre-scripted interaction and conducting the following five tasks sequentially: (i) Task 0: Icebreaking task; (ii) Task 1: Recall of a happy memory and a sad memory; (iii) Task 2: Short Moods and Feelings Questionnaire (SMFQ) [66]; (iv) Task 3: Picture task inspired by the Child Apperception Test (CAT) [67]; and (v) Task 4: three subscales from the Revised Child Anxiety and Depression Scale (RCADS) [68]. Further details regarding the experimental setup and procedure can be found in [4] and[62].
@@
In order to compare whether children’s responses to a robot that is remotely connected via Zoom or Skype differ from the children’s responses to a robot that is co-located, in the context of mental wellbeing assessment, a similar experimental protocol and tasks were adopted for the first session of this study. In this paper, we will refer to the study that used online interactions with the robot as the "remote study" while our previous study undertaken in the lab introduced in [4, 62] will be referred to as the "in-lab study".
4.1 Materials and Methods
4.1.1 Participants
@@
In planning for a repeated measures analysis to detect a medium effect size, a priori power analysis was conducted using G*power software [69], accounting for three groups and three repetitions per participant. Parameters included a medium effect size f=0.25, an $\alpha$ error probability of 0.05, desired power (1 - $\beta$) of 0.80, correlation among repeated measures of r = 0.5, and a nonsphericity correction of $\epsilon$=0.75, indicating a moderate expectation of sphericity violation. The analysis recommended a total sample size of 36 participants to achieve an actual power (1 - $\beta$) of approximately 0.83, with a noncentrality parameter $\lambda$=10.13, critical F-value of 3.51, and degrees of freedom adjusted for the anticipated non-sphericity (numerator df=1.5, denominator df=49.5). This conservative approach ensures robustness against potential violations of sphericity and realistic assumptions about within-subject correlations, adequately powering the study to detect the specified effect size. To accommodate potential dropouts and ensure we have enough data for the supplementary analyses based on participants’ gender, we have increased the sample size by 10%, aiming for at least 20 units in each cell for drawing statistical inferences [70]. Further stratifications of the data beyond the three tertiles, such as those based on gender, are complementary to the main analysis presented in this paper. The final sample size was determined following previous CRI studies yielding similar results with long-term experimental designs [e.g., 71], showing that a final sample size of 40 participants is appropriate.
@@
42 individuals between the ages of 8 and 13 years old were recruited using one of the two recruitment approaches: (a) advertising using online flyers that were circulated across the schools, and (b) snowball sampling by leveraging contacts among the research team. The recruitment criteria involved exclusion on the basis of any existing neurological or psychological conditions declared by the parents/guardians. Two participants were not included in the study mentioned here: one due to technical difficulties and another because of the inability to participate in the experiment sessions. Thus, the final sample consisted of 40 participants (20 girls and 20 boys, mean age = $10.04$, SD= $1.58$). One participant turned 14 during the course of the study sessions, however, was included in the study sample as they were 13 at the time of recruitment. The participants were from different parts of the world (Tajikistan: 13, UK: 8, Singapore: 7, Qatar: 6, India: 3, Other: 3) and interacted with the robot remotely using Zoom or Skype. Informed written consent was obtained from the parents/guardians prior to the participation of their children. The study was approved by the Cambridge Psychology Research Ethics Committee at the University of Cambridge.
4.1.2 Setup
@@ 4.1.3 Robotic platform
The study was conducted using the Nao humanoid platform from Aldebaran Robotics [^6] https://www.aldebaran.com/en/nao. The Nao robot is equipped with human-like movement, visual sensors for object identification, and speech generation and is quite prevalent in CRI literature [72, 73]. The robot followed a ‘Wizard of Oz’ paradigm in which the interaction dialogues were pre-programmed and did not change with the participants. However, the experiment flow, i.e. the continuation of the activities as to when the robot should respond or proceed with the subsequent activity, was determined by the researcher. This was done so that any interruptions to the experiment session due to the low quality of the conference call or internet issues do not affect the experiment tasks. Only one wizard who was a researcher conducted all the experiment sessions for all the participants.
4.1.4 Experiment protocol
@@
The researchers sent out the link for the video conference call prior to the experiment sessions. The sessions’ dates and times were determined based on the children’s availability. The children interacted with the robot for three sessions at least one week apart. For every interaction session, each child was requested to log in to their preferred video conferencing platform (Zoom or Skype), prior to the experiment session. The experimenter also joined the call using the camera and audio from their laptop and welcomed the child to the session. The experimenter described certain housekeeping rules in relation to the experiment session, including asking the child to speak clearly and sufficiently loudly and telling them about the regular breaks in between the activities. They were also instructed that they could pause the interaction, ask the robot to repeat what it was saying, and stop or skip parts of the activities, in accordance with how they felt. Once the child was comfortable and ready to begin the interaction, the experimenter switched the camera of Zoom or Skype to the Logitech camera placed in front of the robot. This was done so that the experimenters were no longer in the frame of view of the participants. In some cases, the experimenter might also switch the audio input from the laptop computer to the Jabra microphone placed in front of the robot, depending on what the participant felt comfortable with in terms of the audio quality of the conference call.
@@
The first interaction session started with the robot introducing itself, waving to the child, telling some jokes and playing tic tac toe with the child. The second interaction session began with the robot giving a brief introduction again about the session, waving to the child, and telling the children some jokes. Finally, the third interaction session started with the robot introducing the session, waving and playing a guessing game with the children. For all the sessions, the initial introduction task ended with the robot asking the children ice-breaking questions like "how was your day?", in order to get the children comfortable with the robot and its capabilities. After the introduction activity, the robot proceeded to perform the tasks as described in the next section (Sec. 4.1.5 Study tasks). The activity flow was controlled by the experimenter who was responsible for when the robot should proceed with the task. This was determined in accordance with when the child finished with their responses so that interruptions due to internet connection or the quality of the conference call did not affect the completion of the experiment tasks.
4.1.5 Study tasks
@@
In order to investigate whether co-location with the robot affects wellbeing assessment in children, we followed the experiment tasks of our prior work [4] which consisted of measures that are established standards in HRI and child psychology [74, 75, 76]. The first session consisted of the following tasks:
@@
Task2: The robot administered the Short Moods and Feelings Questionnaire (SMFQ) [77]. The robot said aloud statements like "You felt lonely" from the questionnaire and the child could respond by verbalising their responses. The display screen was used to project the response ratings "Not true", "Sometimes" and "True" and the child was instructed by the robot to pick one of the response ratings in accordance with how they were feeling in the last two weeks.
@@
Task3: The robot conducted a picture-based task which was inspired by the Child Apperception Test (CAT) [67]. The robot showed picture 7 from the CAT in the first session, using the computer screen and helped the children to form a narrative by asking questions like "What do you think is happening in this picture?", "What do you think happened before this picture?" and "What do you think happened after this picture?". The task was used to determine children’s wellbeing based on the children’s description of the pictures and the narrative they came up with. Similar tasks have also been used previously in HRI [75].
@@
Task 4: The robot administered three subscales (generalised anxiety, panic disorder and low mood) from the Revised Child Anxiety and Depression Scale (RCADS) [68]. The robot made statements in accordance with the items of the questionnaire like "When you have a problem, your heart beats really fast". The children were instructed to respond using one of the response ratings "Never", "Sometimes", "Often", and "Always" that were displayed on the computer screen.
@@
For study sessions 2 and 3, the order of the tasks was changed in the following manner: (1) Task 1- recalling a recent happy and sad memory, (2) Task 2- the robot conducted the RCADS, and (3) Task 3- the robot administered the picture based task (picture 9 from the CAT in session 2 and picture 10 from the CAT in session 3 [67]). The SMFQ task was only present in the first session. The experiment sessions concluded with the robot either playing a game with the child (session 1 and session 2) or telling some jokes to the child (session 3). All three sessions concluded with the experimenter conducting the user perception questionnaires. Since one of the aims of this work is to understand whether co-location affects the perception of the robot for mental wellbeing assessment, we adopted the user perception questionnaires employed in our prior work mentioned in [62]. The first questionnaire was inspired by the questionnaire in [78] to measure the children’s perception of the robot as a confidante. Open-ended questions and statement 10 ("Talking with the robot was similar to talking with a stranger") were excluded. The children could respond using Likert ratings: (1) Strongly dislike, (2) Somewhat dislike, (3) Neither like nor dislike, (4) Somewhat like, and (5) Strongly like. Children’s anxiety towards robots was measured using a questionnaire inspired by the RAS [79]. To enable clarity and better understanding for children, statement 3 of RAS was modified to "The robot may be unable to understand my stories", statement 10 was changed to "Whether the robot understands what I am saying", and statement 11 was modified to "I may be unable to understand what the robot is saying". Children could respond by verbalising one of the response ratings: (1) I do not feel anxiety at all, (2) I hardly feel any anxiety, (3) I don’t feel much anxiety, (4) I feel a little anxiety, (5) I feel much anxiety, and (6) I feel anxiety very strongly. Finally, the third user perception questionnaire was inspired by the Godspeed questionnaire [80]. We used the likeability and perceived intelligence subsections of the questionnaire and the children could respond using the questionnaire response ratings ranging from one to five (one corresponded with the negative characteristic like "ignorant" and five corresponded with the positive characteristic like "knowledgeable").
4.1.6 Measures

### v3

@@ 2.7. Research Questions and Contributions
Children differ widely in their attitudes towards technology which might affect their interaction with a social robot (Jackson et al., 2008). This is of specific importance when encouraging them to open up about their wellbeing. It has also been shown that wellbeing is influenced by multiple factors (family problems, financial issues, academic workload etc.) and is subject to change without any warning or discerning pattern (Gromada et al., 2020). Thus, we investigated whether and how online interaction (using Zoom or Skype) with a social robot can be used for the longitudinal assessment (three sessions over the course of 3-6 weeks) of mental wellbeing (RQ1). Longitudinal CRI studies have also observed that habituation with the robot leads to children losing interest in the interaction, causing less engagement and attention to tasks (Salter et al., 2004). Thus, we investigated how children’s perception of the robot changed over the course of the longitudinal interaction (RQ2). Specifically, we also investigated whether remote interactions with the robot affect the children’s perception of the robot as compared to co-location (RQ2.1). Gender-based differences have also been observed with regard to participants’ attitudes towards mental health (Chandra and Minkovitz, 2006) and their HRI experience (Strait et al., 2015). Thus, we also conducted a supplementary exploratory investigation into whether and how gender affects the robot-assisted longitudinal assessment of the mental wellbeing of children and their perception of the robot’s behaviour and capabilities (RQ3).
@@
Figure 2. (a) Researcher view of the experimental setup, (b) Participant view of the interaction.
3. Methods
The following section describes the study design, participants, experimental protocol and data analyses conducted in this work.
3.1. In-lab Study
@@
The experiment design for the present work was motivated by our prior investigation examining whether and how a humanoid robot can be used for the evaluation of mental wellbeing of children in a dyadic physical lab setting (Abbasi et al., 2022, 2023, 2024b). In our in-lab study, we collected data from 41 children [^2] In order to balance out the gender of the participants, we collected data from about 13 more participants (as mentioned in (Abbasi et al., 2024b)) after the 28 participants mentioned in (Abbasi et al., 2022) (21 girls, 20 boys, mean age = $9.58y.o.$, SD = $1.45y.o.$). The participants were recruited from Cambridgeshire, United Kingdom. The experimental protocol consisted of the robot following a pre-scripted interaction and conducting the following five tasks sequentially: (i) Task 0: Icebreaking task; (ii) Task 1: Recall of a happy memory and a sad memory; (iii) Task 2: Short Moods and Feelings Questionnaire (SMFQ) (Angold et al., 1995); (iv) Task 3: Picture task inspired by the Child Apperception Test (CAT) (Bellak and Bellak, 1949); and (v) Task 4: three subscales from the Revised Child Anxiety and Depression Scale (RCADS) (Chorpita et al., 2015). Further details regarding the experimental setup and procedure can be found in (Abbasi et al., 2022),(Abbasi et al., 2023) and (Abbasi et al., 2024b).
@@
In order to compare whether children’s responses to a robot that is remotely connected via Zoom or Skype differ from the children’s responses to a robot that is co-located, in the context of mental wellbeing assessment, a similar experimental protocol and tasks were adopted for the first session of this study. In this paper, we will refer to the study that used online interactions with the robot as the ”remote study” while our previous study undertaken in the lab introduced in (Abbasi et al., 2022, 2023, 2024b) will be referred to as the ”in-lab study”. It should be noted that new participants were recruited for the remote study in order to ensure that no individual participated in both the remote study and the in-lab study.
3.2. Remote Study
The experiment protocol followed during the remote study has been mentioned below. It should be noted that in order to make valid comparison with our previous works (Abbasi et al., 2023, 2022, 2024b), we have followed the same experimental protocol in the remote study, only changing the mode of interaction from in-person to online.
3.2.1. Participants
@@
In planning for a repeated measures analysis to detect a medium effect size, a priori power analysis was conducted using G*power software (Faul et al., 2009), accounting for three groups and three repetitions per participant. Parameters included a medium effect size f=0.25, an $\alpha$ error probability of 0.05, desired power (1 - $\beta$) of 0.80, correlation among repeated measures of r = 0.5, and a nonsphericity correction of $\epsilon$=0.75, indicating a moderate expectation of sphericity violation. The analysis recommended a total sample size of 36 participants to achieve an actual power (1 - $\beta$) of approximately 0.83, with a noncentrality parameter $\lambda$=10.13, critical F-value of 3.51, and degrees of freedom adjusted for the anticipated non-sphericity (numerator df=1.5, denominator df=49.5). This conservative approach ensures robustness against potential violations of sphericity and realistic assumptions about within-subject correlations, adequately powering the study to detect the specified effect size. To accommodate potential dropouts and ensure we have enough data for the supplementary analyses based on participants’ gender, we have increased the sample size by 10%, aiming for at least 20 units in each cell for drawing statistical inferences (Simmons et al., 2011). Further stratifications of the data beyond the three tertiles, such as those based on gender, are complementary to the main analysis presented in this paper. The final sample size was determined following previous CRI studies yielding similar results with long-term experimental designs (e.g., Ahmad et al., 2017), showing that a final sample size of 40 participants is appropriate.
@@
42 individuals between the ages of 8 and 13 years old were recruited using one of the two recruitment approaches: (a) advertising using online flyers that were circulated across the schools, and (b) snowball sampling by leveraging contacts among the research team. Informed written consent was obtained from the parents/guardians prior to the participation of their children. The recruitment criteria involved exclusion on the basis of any existing neurological or psychological conditions declared by the parents/guardians. The recruitment criteria also included that the children were fluent in English which was the medium of communication of the study. This was confirmed based on the language of instruction in their respective schools to be English. The parents were asked to verify this on the consent form. Two participants were not included in the study mentioned here: one due to technical difficulties and another because of the inability to participate in the experiment sessions. Thus, the final sample consisted of 40 participants (20 girls and 20 boys, mean age = $10.04$, SD= $1.58$). One participant turned 14 during the course of the study sessions, however, was included in the study sample as they were 13 at the time of recruitment. The participants were from different parts of the world (Tajikistan: 13, UK: 8, Singapore: 7, Qatar: 6, India: 3, Other: 3) and interacted with the robot remotely using Zoom or Skype. The study was approved by the Cambridge Psychology Research Ethics Committee at the University of Cambridge.
3.2.2. Setup
@@ 3.2.3. Robotic platform
The study was conducted using the Nao humanoid platform from Aldebaran Robotics [^3] https://www.aldebaran.com/en/nao. The Nao robot is equipped with human-like movement, visual sensors for object identification, and speech generation and is quite prevalent in CRI literature (de Greeff et al., 2014; Arsić et al., 2022). The robot followed a ‘Wizard of Oz’ paradigm in which the interaction dialogues were pre-programmed and did not change with the participants. However, the experiment flow, i.e. the continuation of the activities as to when the robot should respond or proceed with the subsequent activity, was determined by the researcher. This was done so that any interruptions to the experiment session due to the low quality of the conference call or internet issues do not affect the experiment tasks. Only one wizard who was a researcher conducted all the experiment sessions for all the participants.
3.2.4. Experiment protocol
@@
The researchers sent out the link for the video conference call prior to the experiment sessions. The sessions’ dates and times were determined based on the children’s availability. The children interacted with the robot for three sessions at least one week apart. For every interaction session, each child was requested to log in to their preferred video conferencing platform (Zoom or Skype), prior to the experiment session. The experimenter also joined the call using the camera and audio from their laptop and welcomed the child to the session. The experimenter described certain housekeeping rules in relation to the experiment session, including asking the child to speak clearly and sufficiently loudly and telling them about the regular breaks in between the activities. They were also instructed that they could pause the interaction, ask the robot to repeat what it was saying, and stop or skip parts of the activities, in accordance with how they felt. Once the child was comfortable and ready to begin the interaction, the experimenter switched the camera of Zoom or Skype to the Logitech camera placed in front of the robot. This was done so that the experimenters were no longer in the frame of view of the participants. In some cases, the experimenter might also switch the audio input from the laptop computer to the Jabra microphone placed in front of the robot, depending on what the participant felt comfortable with in terms of the audio quality of the conference call. The following are the steps that were taken to ensure that ethical practices during the child-robot interaction sessions:
  1. (1)
    Communication of study procedures: The experimenters made efforts to ensure that the study procedures were communicated with children in an age-appropriate manner. This was done verbally at the start of the sessions as well as throughout the session. The experimenters used simple and non-technical language to explain the experimental sessions and the role of the robot without delving into the study tasks in order to avoid priming and preparedness for answers. They ensured the children understood that they were not required to perform any task they did not want to.
  2. (2)
    Study tasks and breaks: After each study task, the robot would ask the participants to see if they wanted to continue or take a break. If the children decided to take a break, the experimenter would come back into the field of view of the children and ensure that they felt comfortable resuming the study when they were ready to do so. The practice of allowing the children to take breaks and check in was designed in order to maintain a comfortable and supportive environment for the children.
  3. (3)
    Informed consent and assent: In addition to obtaining written consent from the parents, the study employed a child assent process to ensure that the children were comfortable with their participation. The experimenters verbally reassured that they could stop the session at any time. This opportunity to withdraw was clearly communicated to the children and they were reminded that they could take breaks and discontinue the interaction at any point they wished.
  4. (4)
    Ethical practices for child-robot interaction: Ethical considerations were at the forefront of the study design. Several measures were taken to minimise any risk of harm:
    1. (a)
      Safety of the robot: The robot was tested and designed with the children’s safety in mind. It was ensured that the robot was safe to interact with and any potential for distress was minimised by pre-testing the interactions.
    2. (b)
      Supervision and monitoring: All interactions were monitored by research staff to address any concerns immediately. The children were encouraged to express their feelings and any signs of discomfort were addressed promptly.
    3. (c)
      Informed parents: Parents were informed about the nature of the study prior to the experiment sessions through the participant information sheet. They had the option to contact the research staff for any clarification/copy of the study tasks in order to ensure transparency and ongoing consent. They were also informed that they could contact the research staff if they would like to discuss the responses of their wards.
  5. (5)
    Recording and privacy: The children were made aware that they were being recorded and they had the option to end the call as mentioned previously.
@@
The first interaction session started with the robot introducing itself, waving to the child, telling some jokes and playing tic tac toe with the child. The second interaction session began with the robot giving a brief introduction again about the session, waving to the child, and telling the children some jokes. Finally, the third interaction session started with the robot introducing the session, waving and playing a guessing game with the children. The decision to conduct three sessions was informed by prior literature (Nalin et al., 2012a; Ahmad et al., 2017) to capture initial reactions as well as allow for a more nuanced understanding of how children would interact with the robot differently over time. For all the sessions, the initial introduction task ended with the robot asking the children ice-breaking questions like ”how was your day?”, in order to get the children comfortable with the robot and its capabilities. After the introduction activity, the robot proceeded to perform the tasks as described in the next section (Sec. 3.2.5 Study tasks). The activity flow was controlled by the experimenter who was responsible for when the robot should proceed with the task. This was determined in accordance with when the child finished with their responses so that interruptions due to internet connection or the quality of the conference call did not affect the completion of the experiment tasks.
3.2.5. Study tasks
@@
In order to investigate whether co-location with the robot affects wellbeing assessment in children, we followed the experiment tasks of our prior work (Abbasi et al., 2022, 2024b) which consisted of measures that are established standards in HRI and child psychology (Cervin et al., 2022; Bremner et al., 2016; Di Nuovo et al., 2019). The first session consisted of the following tasks:
@@
Task2: The robot administered the Short Moods and Feelings Questionnaire (SMFQ) (Angold et al., 1995; Sharp et al., 2006). The robot said aloud statements like ”You felt lonely” from the questionnaire and the child could respond by verbalising their responses. The display screen was used to project the response ratings ”Not true”, ”Sometimes” and ”True” and the child was instructed by the robot to pick one of the response ratings in accordance with how they were feeling in the last two weeks.
@@
Task3: The robot conducted a picture-based task which was inspired by the Child Apperception Test (CAT) (Bellak and Bellak, 1949). The robot showed picture 7 from the CAT in the first session, using the computer screen and helped the children to form a narrative by asking questions like ”What do you think is happening in this picture?”, ”What do you think happened before this picture?” and ”What do you think happened after this picture?”. The task was used to determine children’s wellbeing based on the children’s description of the pictures and the narrative they came up with. Similar tasks have also been used previously in HRI (Bremner et al., 2016).
@@
Task 4: The robot administered three subscales (generalised anxiety, panic disorder and low mood) from the Revised Child Anxiety and Depression Scale (RCADS) (Chorpita et al., 2015). The robot made statements in accordance with the items of the questionnaire like ”When you have a problem, your heart beats really fast”. The children were instructed to respond using one of the response ratings ”Never”, ”Sometimes”, ”Often”, and ”Always” that were displayed on the computer screen.
@@
For study sessions 2 and 3, the order of the tasks was changed in the following manner: (1) Task 1- recalling a recent happy and sad memory, (2) Task 2- the robot conducted the RCADS, and (3) Task 3- the robot administered the picture based task (picture 9 from the CAT in session 2 and picture 10 from the CAT in session 3 (Bellak and Bellak, 1949)). The SMFQ task was only present in the first session. The experiment sessions concluded with the robot either playing a game with the child (session 1 and session 2) or telling some jokes to the child (session 3). All three sessions concluded with the experimenter conducting the user perception questionnaires. Since one of the aims of this work is to understand whether co-location affects the perception of the robot for mental wellbeing assessment, we adopted the user perception questionnaires employed in our prior work mentioned in (Abbasi et al., 2023). The first questionnaire was inspired by the questionnaire in (Murray et al., 2022) to measure the children’s perception of the robot as a confidante. Open-ended questions and statement 10 (”Talking with the robot was similar to talking with a stranger”) were excluded. The children could respond using Likert ratings: (1) Strongly dislike, (2) Somewhat dislike, (3) Neither like nor dislike, (4) Somewhat like, and (5) Strongly like. Children’s anxiety towards robots was measured using a questionnaire inspired by the RAS (Nomura et al., 2006). To enable clarity and better understanding for children, statement 3 of RAS was modified to ”The robot may be unable to understand my stories”, statement 10 was changed to ”Whether the robot understands what I am saying”, and statement 11 was modified to ”I may be unable to understand what the robot is saying”. Children could respond by verbalising one of the response ratings: (1) I do not feel anxiety at all, (2) I hardly feel any anxiety, (3) I don’t feel much anxiety, (4) I feel a little anxiety, (5) I feel much anxiety, and (6) I feel anxiety very strongly. Finally, the third user perception questionnaire was inspired by the Godspeed questionnaire (Bartneck et al., 2009). We used the likeability and perceived intelligence subsections of the questionnaire and the children could respond using the questionnaire response ratings ranging from one to five (one corresponded with the negative characteristic like ”ignorant” and five corresponded with the positive characteristic like ”knowledgeable”). It should be noted that in order to maintain a uniform interaction experience for all participants, the robot behaviour was consistent. In other words, the robot did not build off the prior interactions with the participants.
3.2.6. Measures